OpenStudy (anonymous):

Will fan and medal... Identify the sequence graphed below and the average rate of change from n = 1 to n = 3.

OpenStudy (anonymous):

The graph shows the points.. (2,8) (3,4) (4,2) (5,1)

OpenStudy (anonymous):

OK. I really don't know what they would want from you when they say "identify" the sequence, but (x,y) --> (x+1,y/2) is the pattern that I see.

OpenStudy (anonymous):

Yea, figured out the 1/2 part of the equation, but i need the entire equation and the average rate of change...

OpenStudy (dan815):

hey

OpenStudy (anonymous):

hey. can u help me?

OpenStudy (dan815):

yes

OpenStudy (dan815):

so we are doing divisions by 2

OpenStudy (dan815):

y=8/2^(x-2)

OpenStudy (anonymous):

so 8 would be the a1 value???

OpenStudy (dan815):

that is the term 1 yah, what you mean by a1

OpenStudy (dan815):

rate*

OpenStudy (dan815):

divided by 2* not 3

OpenStudy (anonymous):

im so confused. the options for the rate of change are 6 and -6

OpenStudy (dan815):

its (2-8)/2 =-6/2 = -3

OpenStudy (dan815):

hmm how do they get -6

OpenStudy (dan815):

go with first choice

OpenStudy (anonymous):

ok thanks. do you think you can help with one more?

OpenStudy (dan815):

ok i will try

OpenStudy (dan815):

:)

OpenStudy (anonymous):

(cos Θ − cos Θ)^2 + (cos Θ + cos Θ)^2

OpenStudy (dan815):

expand them and add them it will simplify

OpenStudy (dan815):

are you sure thats how it is

OpenStudy (dan815):

why arent they just subtracting the left side its 0^2

OpenStudy (dan815):

and you have 4cos^2(theta)

OpenStudy (anonymous):

thank you :)

OpenStudy (dan815):

you're welcome see ya later!